how many solutions does the linear system 3x + 5y = 8 and 3x + 5y = 1 have?
Answer by stanbon(75887)

how many solutions does the linear system 3x + 5y = 8 and 3x + 5y = 1 have?
----------
solve each equation for y:
y = (8-3x)/5; slope = -3/5
y = (1-3x)/5; slope = -3/5
-----------------------------
The lines are parallel so there is no solution for the system.
================================
Cheers,
Stan H.